OpenStudy (anonymous):

39.5454545, rounded to the nearest tenth ?

OpenStudy (anonymous):

Find the tenths place, look at the next number to the right. If the number is five or greater then round up. If the number is less than five round down.

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

For example : 9.4 we cant say = 9+1 but : 9.5 can say = 9+1

OpenStudy (anonymous):

want me to help or u got it?

OpenStudy (anonymous):

i got it , thanks (: